OpenStudy (anonymous):

Suppose you solved a system of linear equations using substitution and you got the results given. How many solutions does each system have? -7=4 is no solution correct?

OpenStudy (anonymous):

for linear function, the solution should only one; when you try to solve the problem, try to get rid of x or y

OpenStudy (aroub):

correct!! :)

OpenStudy (aroub):

there is no solution , FASLE!

OpenStudy (anonymous):

so im correct

OpenStudy (anonymous):

there is 1

OpenStudy (anonymous):

??

OpenStudy (anonymous):

ex 1: x-y=1, 3x-6y=10, 3x-3y=3,3x-6y=10, 3y=-7

OpenStudy (aroub):

idk u just cant have -7=5 its false!!

OpenStudy (anonymous):

so its no solution, or 1 solution

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i still think none

OpenStudy (anonymous):

actually. no answer, one answer, or many answers

OpenStudy (anonymous):

?? which 1??

OpenStudy (anonymous):

0??

OpenStudy (anonymous):

cuz 4 cant = -7

OpenStudy (anonymous):

maybe I think too much, just forget it

OpenStudy (anonymous):

so wat is it?? do u think

OpenStudy (anonymous):

for this question ,no answer

OpenStudy (anonymous):

and if its x=-1 that is 1 solution?

OpenStudy (anonymous):

yes, x=-1, y=3,4,5,....

OpenStudy (anonymous):

wat happens if the numbers are = like 9=9

OpenStudy (anonymous):

please answer :)

OpenStudy (anonymous):

I think this is only a point :9

OpenStudy (anonymous):

so it 1 correct

OpenStudy (anonymous):

yes